OpenStudy (anonymous):

how do you factor 7x - 60 + x^2? please show all steps to get to the answer

OpenStudy (anonymous):

STEP1: Solve quadratic equation x2 + 7x - 60 = 0 The solutions are: x1 = 5 x2 = -12 STEP2: Use formula for factoring quadratic equation: ax2 + bx + c = a(x - x1)(x - x2) Where x1 and x2 are solutions of quadratic equation In our case we have: x2 + 7x - 60 = 1(x - 5)(x - (-12)) that is x2 + 7x - 60 = (x - 5)(x + 12) answer: (x - 5)(x + 12) I hope I helped.

OpenStudy (anonymous):

thanks so much!! you did help a lot :) @arianadarsh

OpenStudy (anonymous):

your welcome any time